# Smoothly vanishing density in the contact process by an interplay of disorder and long-distance dispersal

### v1

@@ I Introduction
This simple CP has been extended in at least two directions with the purpose of a more realistic modeling of ecological systems. First, according to observations, colonization events occur also to distant places, and the dispersal is heavy-tailed, i.e., the colonization probability decreases with the distance slower-than-exponentially [8, 9, 10, 11, 12, 13, 14]. This issue has been studied in the CP for the case of a dispersal probability decaying algebraically with the distance as $p(l)\sim l^{-\alpha}$ [15, 16, 17]. According to these studies, the critical behavior remains in the DP class if $\alpha>\alpha_{c}(d)$, where $\alpha_{c}(d)$ is a dimension-dependent threshold value, while, for $\alpha<\alpha_{c}(d)$, the critical exponents are altered and depend on $\alpha$. Thus, even in this case, the algebraic vanishing of the order parameter in Eq. (1) holds to be valid, although with a modified $\beta$. Second, real systems may have a fine-scale heterogeneity in the conditions of surviving or reproduction. This can be taken into account in the CP by quenched random colonization and extinction rates. The well known, perturbative Harris criterion [18] predicts weak disorder to be relevant in low dimensions ($d<4$), leading to a new type of critical behavior. According to the strong-disorder renormalization group (SDRG) method [19, 20], at least for sufficiently strong disorder [21], the critical behavior is described by an infinitely disordered fixed point (IDFP) in agreement with results of Monte Carlo simulations [24, 25, 26]. At the IDFP, the dynamics is highly uncommon, as time itself is replaced by its logarithm in scaling relations, but otherwise power-laws like Eq. (1) hold to be valid with critical exponents different from those of DP.
@@
with $C$ being a nonuniversal positive constant. This type of infinite-order vanishing of the density, corresponding formally to $\beta=\infty$, is qualitatively different from the conventional power-law singularity in Eq. (1).
@@ II.2 The SDRG method
When these operations are attempted in $d>1$ dimensions, one encounters a difficulty of arising double activation rates between clusters. We treated this problem by following the standard ‘maximum rule’, which means that only the larger one of the rates is kept. A further modification in the rule in Eq. (6) we have done is that $\kappa=1$ has been used instead of $\kappa=2$. The reason for this, is that this modified rule, together with the maximum rule allows us to use an efficient numerical algorithm of the SDRG procedure developed by one of us [39]. For the CP with nearest-neighbor dispersal on low-dimensional lattices, where the critical fixed point of the SDRG transformation is an IDFP, both the maximum rule and the irrelevance of the value of $\kappa$ are justified [40]. For the CP with long-distance (algebraic) dispersal, where the critical fixed point is an FDFP [27], Monte Carlo results on the dynamical scaling at the critical point in $d=1$ and $d=2$ confirmed the validity of these approximations for $\alpha>\frac{3}{2}d$, and also at $\alpha=\frac{3}{2}d$ for a low enough dilution [28, 30].
@@ II.3 Monte Carlo simulation
In the simulations, a different type of disorder, which is more frequently used in Monte Carlo studies [25, 26, 28], was considered. This is a random dilution of the lattice, by which a fraction $c$ of sites is randomly deleted. Note that, due to long-distance dispersal, $c$ does not need to be below the percolation threshold of the underlying regular lattice. We applied a sequential update, in which an active site is picked randomly and either made inactive with a probability $1/(1+\lambda)$ or an activation event is attempted from this site with a probability $\lambda/(1+\lambda)$. In the latter case, a target site is selected randomly with a probability proportional to $l^{-\alpha}$, where $l$ is the Euclidean distance measured from the source site. As control parameter of the transition we used $\lambda$. Technical details of the simulation are the same as in Ref. [28].
@@ III.1 The one-dimensional model
For the one-dimensional model, a simplified, analytically tractable SDRG scheme predicts a smooth vanishing of the global density in the active phase as given by Eq. (2) [31]. To check the validity of this asymptotic form, we performed Monte Carlo simulations for $\alpha=2$ and $\alpha=3$, with a dilution parameter $c=0.5$. Data obtained for different values of the control parameter $\lambda$ and at different times are plotted in Fig. 1. Plotting the logarithmic densities $\ln\rho(\lambda,t)$ against $\Delta^{-1/2}$, with $\Delta=\lambda-\lambda_{c}$, leads to an agreement with Eq. (2), as the late-time densities (which approached their steady-state values) fit well to a straight line. Here, estimates of the critical control parameter $\lambda_{c}$ were taken from Ref. [28]. We find the constant $C$ appearing in Eq. (2) to vary with $\alpha$.
@@ III.2 The two-dimensional model
Refer to captionRefer to caption
Figure 6: Dependence of the persistence probability on the control parameter for different system sizes. Data have been obtained by Monte Carlo simulations of the two-dimensional model with $\alpha=3$ (a) and $\alpha=3.5$ (b). The straight lines have slopes $\beta_{p}^{\prime}=0.67$ (a) and $\beta_{p}^{\prime}=0.69$ (b).
These confirm what has been found by the SDRG method: the vanishing is of mixed-order type. The estimated critical persistences are $\pi_{0}=0.042(2)$ for $\alpha=3$ and $\pi_{0}=0.070(2)$ for $\alpha=3.5$. Since the value of $\beta_{p}^{\prime}$ obtained by fitting is very sensitive to the errors of $\lambda_{c}$ and $\pi_{0}$, we have rather uncertain estimates on it: $\beta_{p}^{\prime}=0.7(2)$ both for $\alpha=3$ and $\alpha=3.5$. Note that, just like the two-dimensional SDRG results, these do not exclude a possible agreement with the value $1/2$ obtained by the SDRG method for $d=1$.
@@ IV Discussion
In this work, we considered a stochastic lattice model of population dynamics, the contact process in the simultaneous presence of quenched disorder and long-distance dispersal. We focused on the off-critical scaling of order parameters, for which an earlier SDRG work predicted anomalous scaling behavior in one dimension, for large enough dispersal exponents. Specifically, the global density vanishes smoothly, whereas the persistence probability vanishes discontinuously when the critical point is approached from the active phase and the inactive phase, respectively [31]. We provided a numerical confirmation of these forms by Monte Carlo simulations in one dimension and, making a step toward a more realistic modelling, we addressed this issue in the two-dimensional model. In two dimensions, our numerical SDRG and Monte Carlo results agree in that the global density follows the same type of smooth vanishing given in Eq. (2) as found in one dimension. Furthermore, the discontinuous, mixed-order type of vanishing of the persistence found in one dimension holds to be valid also in two dimensions, as it is demonstrated by our numerical SDRG and Monte Carlo results. Concerning the exponent $\beta_{p}^{\prime}$ in the subleading term of persistence, our numerical results clearly signal a singularity with $\beta_{p}^{\prime}<1$, but the accessible numerical data are not conclusive on whether this exponent is different from that of the one-dimensional model ($1/2$) or not.
@@
There are several directions in which the investigations performed in this work could be extended. From a theoretical perspective, one may pose the question whether the anomalous behavior of order parameters found for $d=1$ and $d=2$ holds to be valid in higher dimensions, as well. Based on the conjectured connection between different dimensional variants of the model, we expect qualitatively similar results in three dimensions (with a possibly different value of $\beta_{p}^{\prime}$) in the non-mean-field domain $\alpha>\frac{9}{2}$, where the Harris criterion predicts weak disorder to be relevant. However, in four dimensions and above, where weak disorder is irrelevant for any value of the dispersal exponent $\alpha$, we expect the model to exhibit mean-field critical behavior. With the scope of ecological modeling, it would be interesting to analyse the model under an environmental gradient [41]. This is naively expected to transform the control-parameter dependence of global density examined in this paper to a spatial dependence of the local density along the direction of the gradient. Moreover, this would allow us to confront the predictions of the model with observational data obtained by satellite images in the presence of an environmental gradient.

### v2

@@ I Introduction
This simple CP has been extended in at least two directions with the purpose of a more realistic modeling of ecological systems. First, according to observations, colonization events occur also to distant places, and the dispersal is heavy-tailed, i.e., the colonization probability decreases with the distance slower-than-exponentially [8, 9, 10, 11, 12, 13, 14]. This issue has been studied in the CP for the case of a dispersal probability decaying algebraically with the distance as $p(l)\sim l^{-\alpha}$ [15, 16, 17]. According to these studies, in low dimensions ($d<4$), the critical behavior remains in the DP class if $\alpha>\alpha_{c}(d)$, where $\alpha_{c}(d)$ is a dimension-dependent threshold value, while, for $\alpha_{\rm MF}(d)<\alpha<\alpha_{c}(d)$, the critical exponents are altered and depend on $\alpha$. Finally, for $\alpha<\alpha_{\rm MF}(d)=\frac{3}{2}d$, the system enters the mean-field regime. Thus, even in this case, the algebraic vanishing of the order parameter in Eq. (1) holds to be valid, although with a modified $\beta$. Second, real systems may have a fine-scale heterogeneity in the conditions of surviving or reproduction. This can be taken into account in the CP by quenched random colonization and extinction rates. The well known, perturbative Harris criterion [18] predicts weak disorder to be relevant in low dimensions ($d<4$), leading to a new type of critical behavior. According to the strong-disorder renormalization group (SDRG) method [19, 20], at least for sufficiently strong disorder [21], the critical behavior is described by an infinitely disordered fixed point (IDFP) in agreement with results of Monte Carlo simulations [24, 25, 26]. At the IDFP, the dynamics is highly uncommon, as time itself is replaced by its logarithm in scaling relations, but otherwise power-laws like Eq. (1) hold to be valid with critical exponents different from those of DP.
@@
with $C$ being a nonuniversal positive constant. This type of infinite-order vanishing of the density, corresponding formally to $\beta=\infty$, is different from the conventional power-law singularity in Eq. (1).
@@ II.2 The SDRG method
When these operations are attempted in $d>1$ dimensions, one encounters a difficulty of arising double activation rates between clusters. We treated this problem by following the standard ‘maximum rule’, which means that only the larger one of the rates is kept. A further modification in the rule in Eq. (6) we have done is that $\kappa=1$ has been used instead of $\kappa=2$. The reason for this, is that this modified rule, together with the maximum rule allows us to use an efficient numerical algorithm of the SDRG procedure developed by one of us [39]. This implementation of the SDRG method works in a parallel manner, relying on graph algorithms to obtain equivalent results to the above mentioned conceptual picture, but in nearly linear time as a function of the number of sites. For the CP with nearest-neighbor dispersal on low-dimensional lattices, where the critical fixed point of the SDRG transformation is an IDFP, both the maximum rule and the irrelevance of the value of $\kappa$ are justified [40]. For the CP with long-distance (algebraic) dispersal, where the critical fixed point is an FDFP [27], Monte Carlo results on the dynamical scaling at the critical point in $d=1$ and $d=2$ confirmed the validity of these approximations for $\alpha>\frac{3}{2}d$, and also at $\alpha=\frac{3}{2}d$ for a high enough dilution [28, 30].
@@ II.3 Monte Carlo simulation
In the simulations, a different type of disorder, which is more frequently used in Monte Carlo studies [25, 26, 28], was considered. This is a random dilution of the lattice, by which a fraction $c$ of sites is randomly deleted. For practical reasons, it is expedient to use high values of $c$, as low dilutions induce an undesirable crossover from the critical behavior of the clean system, shrinking the true asymptotic behavior to the close vicinity of the critical point. Note that, due to long-distance dispersal, $c$ does not even need to be below the percolation threshold of the underlying regular lattice.
We applied a sequential update, in which an active site is picked randomly and either made inactive with a probability $1/(1+\lambda)$ or an activation event is attempted from this site with a probability $\lambda/(1+\lambda)$. In the latter case, a target site is selected randomly with a probability proportional to $l^{-\alpha}$, where $l$ is the Euclidean distance measured from the source site. As control parameter of the transition we used $\lambda$. Technical details of the simulation are the same as in Ref. [28].
@@ III.1 The one-dimensional model
The one-dimensional model has been studied by a simplified, analytically tractable SDRG scheme which contained a series of approximations [27, 31]. First, at any stage of the renormalization, the only interactions that were taken into account were those between adjacent clusters. Furthermore, these couplings were approximated by the long-range interaction between the closest constituent sites. Up to this point, the scheme is essentially equivalent to the application of the maximum rule. Second, the extension of clusters was neglected compared to the spacings between them. This step is justified in the inactive phase and in the active phase close to the critical point. This simplified method predicted a smooth vanishing of the global density in the active phase as given by Eq. (2) [31]. To check the validity of this asymptotic form, we performed Monte Carlo simulations for $\alpha=2$ and $\alpha=3$, with a dilution parameter $c=0.5$. Data obtained for different values of the control parameter $\lambda$ and at different times are plotted in Fig. 1. Plotting the logarithmic densities $\ln\rho(\lambda,t)$ against $\Delta^{-1/2}$, with $\Delta=\lambda-\lambda_{c}$, leads to an agreement with Eq. (2), as the late-time densities (which approached their steady-state values) fit well to a straight line. Here, estimates of the critical control parameter $\lambda_{c}$ were taken from Ref. [28]. We find the constant $C$ appearing in Eq. (2) to vary with $\alpha$.
@@ III.2 The two-dimensional model
Figure 6: Dependence of the persistence probability on the control parameter for different system sizes. Data have been obtained by Monte Carlo simulations of the two-dimensional model with $\alpha=3$ (a) and $\alpha=3.5$ (b). The straight lines have slopes $\beta_{p}^{\prime}=0.69$ (a) and $\beta_{p}^{\prime}=0.68$ (b).
These confirm what has been found by the SDRG method: the vanishing is of mixed-order type. The estimated critical persistences are $\pi_{0}=0.043(2)$ for $\alpha=3$ and $\pi_{0}=0.070(2)$ for $\alpha=3.5$. Since the value of $\beta_{p}^{\prime}$ obtained by fitting is very sensitive to the errors of $\lambda_{c}$ and $\pi_{0}$, we have rather uncertain estimates on it: $\beta_{p}^{\prime}=0.7(2)$ both for $\alpha=3$ and $\alpha=3.5$. Note that, just like the two-dimensional SDRG results, these do not exclude a possible agreement with the value $1/2$ obtained by the SDRG method for $d=1$.
@@ IV Discussion
In this work, we considered a stochastic lattice model of population dynamics, the contact process in the simultaneous presence of quenched disorder and long-distance dispersal. We focused on the off-critical scaling of order parameters, for which an earlier SDRG work predicted anomalous scaling behavior in one dimension, for large enough dispersal exponents ($\alpha\geq\frac{3}{2}d$). Specifically, the global density vanishes smoothly, whereas the persistence probability vanishes discontinuously when the critical point is approached from the active phase and the inactive phase, respectively [31]. We provided a numerical confirmation of these forms by Monte Carlo simulations in one dimension and, making a step toward a more realistic modelling, we addressed this issue in the two-dimensional model. In two dimensions, our numerical SDRG and Monte Carlo results agree in that the global density follows the same type of smooth vanishing given in Eq. (2) as found in one dimension. Furthermore, the discontinuous, mixed-order type of vanishing of the persistence found in one dimension holds to be valid also in two dimensions, as it is demonstrated by our numerical SDRG and Monte Carlo results. Concerning the exponent $\beta_{p}^{\prime}$ in the subleading term of persistence, our numerical results clearly signal a singularity with $\beta_{p}^{\prime}<1$, but the accessible numerical data are not conclusive on whether this exponent is different from that of the one-dimensional model ($1/2$) or not.
@@
There are several directions in which the investigations performed in this work could be extended. From a theoretical perspective, one may pose the question whether the anomalous behavior of order parameters found for $d=1$ and $d=2$ holds to be valid in higher dimensions, as well. Based on the conjectured connection between different dimensional variants of the model, we expect qualitatively similar results in three dimensions (with a possibly different value of $\beta_{p}^{\prime}$) in the non-mean-field domain $\alpha>\frac{9}{2}$, where the Harris criterion predicts weak disorder to be relevant. However, in four dimensions and above, where weak disorder is irrelevant for any value of the dispersal exponent $\alpha$, just as for $\alpha<\frac{3}{2}d$ in lower dimensions, we expect the model to exhibit mean-field critical behavior characterised with $\beta_{\rm MF}=1$. With the scope of ecological modeling, it would be interesting to analyse the model under an environmental gradient, which is typically realized for plant species living on a hillside, where the environmental factors become more and more unfavorable with increasing or decreasing altitude [41]. In modeling, this circumstance can be implemented by applying a linear spatial trend in the rates along a given direction. Intuitively, such trends are expected to transform the control-parameter dependence of global density examined in this paper to a spatial dependence of the local density along the direction of the gradient. Moreover, the study of environmental gradients would allow us to confront the predictions of the model with observational data accessible by satellite images on the spatial variation of the local density of a given species near the range margin.